OpenStudy (anonymous):

tan(pi/2-theta)cotxtheta-csc^2theta

OpenStudy (goformit100):

I @goformit100 an OS User, Welcomes you to Open Study. 'Discover. Learn. Earn a Certificate Ask questions, give help, and get certified with 150,000 students from 180 countries. ALL THE BEST.

OpenStudy (kira_yamato):

Is it \[\tan \left( \frac{ π }{ 2 } - \theta \right)\cot (x \theta) - \csc^2(\theta)\]? Anyways what is the question?

OpenStudy (anonymous):

kira yes the eqation you have there andd i need to verify trig identities

OpenStudy (kira_yamato):

What's the trig identity you're trying to verify?

OpenStudy (anonymous):

csc^2 theta

OpenStudy (kira_yamato):

So you wanna prove that the whole chunk there equals csc^2 (theta)?

OpenStudy (anonymous):

no i have to translate it to either csc=1/sin or 1+csc^2 ans slove it from there

OpenStudy (kira_yamato):

So it's the whole chunk there equal to csc^2 (theta), solve for theta?

OpenStudy (anonymous):

correct

OpenStudy (kira_yamato):

But do you have any value for x?

OpenStudy (anonymous):

no it dose not have any value

OpenStudy (kira_yamato):

Then I have some problems trying to solve. Sorry...

OpenStudy (anonymous):

its ok no worries, but than you for trying